a. If 5.93 g of the substance melts at 29C, and the heat of fusion for this substance is 154 J/g, how much potential energy was transferred into the sample? (hint 1kJ = 1000J)

b. If 5.93 g of the substance boils at 204C, and the heat of vaporization for this substance is 3240 J/g, how much potential energy was transferred into the sample? (hint 1kJ = 1000J) *
1
Expert's answer
2020-03-25T12:48:22-0400

a). "Q=mLf"

m=5.93g

Lf=154 J/g

Q=5.93*154=913.22J

b)."Q=mHv"

Hv=3240J/g

Q=5.93*3240 =19213.2J